OpenStudy (anonymous):

Application of differentiation. Find the equation of the normal to the curve y=(x+2)e^x-3x at the point (0,2).

OpenStudy (anonymous):

dy/dx at (0,2) = 0 i don't know what to do next.

OpenStudy (anonymous):

@ startanewww what did you get on your derivative at (0,2)?

OpenStudy (anonymous):

isn't it 0?

OpenStudy (anonymous):

after getting the slope at (0,2) find the normal means perpendicular to the line at (0,2)

OpenStudy (anonymous):

1st find the equation of tangent to the curve...

OpenStudy (anonymous):

slope perpendicular to it is m2= - 1/m1

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

but the slope I found for dy/dx at (0,2) is zero. What can i do?

OpenStudy (anonymous):

then, the slope of the normal will be -1/0 = infinity how can i find the equation of normal?

OpenStudy (anonymous):

Diffrentiate y=(x+2)e^x-3x

OpenStudy (anonymous):

u know Product Rule ?

OpenStudy (anonymous):

u get slope = -1

OpenStudy (anonymous):

SO...Slope of Normal = -1 / m1 = -1/-1 = 1

OpenStudy (anonymous):

Nw just write the equation using points (0,2) m = 1 y-y1 = m(x-x1)

OpenStudy (anonymous):

y'=e^x+(x+2)e^x-3 dy/dx at (0,2) = e^0 + (0+2)e^0 - 3 = 0

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

y=(x+2)e^x-3x y' = (x+2) *1 + e^x *1 - 3

OpenStudy (anonymous):

differentiating e^x gives e^x right?

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

lets say y'=0, can you form a line perpendicular to the curve? try it :D

OpenStudy (anonymous):

first fint an equation of the line tangent to the curve at (0,2)

OpenStudy (anonymous):

(y-2)/(x-0)=0 y=2?

OpenStudy (anonymous):

what about the equation of normal?

OpenStudy (anonymous):

yes correct :D then ?

OpenStudy (anonymous):

what can i do?

OpenStudy (anonymous):

the slope will be infinity....

OpenStudy (anonymous):

what is a line perpendicular to line y=2?

OpenStudy (anonymous):

at (0,2) ?

OpenStudy (anonymous):

will the slope of the normal be infinity ?

OpenStudy (anonymous):



OpenStudy (anonymous):

but -1/0 = infinity?

OpenStudy (anonymous):

x=0 is the line

OpenStudy (anonymous):

i don't know the answer....

OpenStudy (anonymous):

i think you are right

OpenStudy (anonymous):

ah ok theres no answer in the book? :D

OpenStudy (anonymous):

my classmates said the answer was x=0

OpenStudy (anonymous):

hmm the line perpendicular to y=2 is the y axis or the vertical line

OpenStudy (anonymous):

ok so the answer should be x=0 only?

OpenStudy (anonymous):

hmm it can be x=0

OpenStudy (anonymous):

x = 0 to infinity?

OpenStudy (anonymous):

it could be y= 0 to 2, or 2 to infinity, or y= 0 to negative infinity

OpenStudy (anonymous):

?????? how come

OpenStudy (anonymous):

equation of the normal at (0,2)...? should there be only one answer?

OpenStudy (anonymous):

because line tangent to the curve is y=2 or the horizontal line, now the line normal to it is the line perpendicular to y=2,, w/c are ________ ?

OpenStudy (anonymous):

vertical lines

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the vertical line y axis

OpenStudy (anonymous):

yess correct :D

OpenStudy (anonymous):

so what is the final answer? x=0 ?